OpenStudy (anonymous):

Create a quadratic function with two real zeros. Solve a system of non-linear equations with a graph. Solve a system of non-linear equations with a table. Compare linear, quadratic, and exponential growth functions using a table and a graph

OpenStudy (anonymous):

you want one with two real zeros, start with it in factored form, like say \[(x-2)(x-3)=0\] then it will have two real zeros

OpenStudy (anonymous):

Two non-linear equations: y = sin x y = x^2 Graph to find x such that x^2 = sin x at the intersections